MODS and LBC checkout

afternoon analyzed focus sequences and examined focus over field for 3 Red and 1 Blue. Blue looked very flat, Red about a 25mic gradient from chip 3 to chip 1.

around 23:30 MST, LBC is ready to go

evening I edited the focus offsets based on the focus sequences obtained last night.

06:30 RPC error on Blue. After recovery, dofpia. -1mm works fine. Converged and corrected pointing, co-pointing and RB.

06:50 Slew to NGC6946 focus. Run dofpia. Seeing estimates from FPIA are not so good - Blue 1.8". See a "skirt" around Blue pupil, which is what is affecting the FPIA-estimate for seeing. We have seen this before. The features in the pupil are sharp, as in very good seeing.

06:57 Start NGC6946 OB. B-BESSEL and R-BESSEL. I will check image quality - to insure focus offsets OK. First images look very nice.

07:10 Uspec and R-BESSEL are 070855 and 070846. Halo around stars on Uspec - is it spherical? V image guide star (for 072019 V and corr red is 072016 which has guide star with 0.8 arcsec seeing est) looks double. Stopping OB and will recollimate.

07:27 V-BESSEL image (072910) definitely shows a field-dependent PSF where on the edge by chip1 (near the guider) the focus/coll really falls apart!

07:33 Recollimate and run OB from SDT_Uspec... 073352. Guider reports 1.1 arcsec Blue and 1.1 Red. Halo around SDT_Uspec images persists. We'll have to check the filter tomorrow afternoon. Halo is not around V-BESSEL stars.

07:52 Focus on Blue starts to fall apart on right side (by guider). Trackers are sending focus corrections. Guide star starts to look double.

08:06 Modified their OB to do V first (so we can check about this falling-apart focus problem) and then U to check on the halos. (OSUmonitorN6946_U). 08:11 see elongated images on one side and round on other side for both channels... so problem may not be isolated to Blue. What is it?

08:30 Field now at airmass 1.6 (obs made from 1.4 to 1.6)
  • Halos around bright stars appear only with SDT_Uspec filter, not with U-Bessel, B-Bessel or V-Bessel. While I took 180s exposures last night at U-BESSEL, I did not take long exposures at SDT_Uspec.
  • Defocus was seen on LBC-Red also - up/down elongated images on left and round images on right. But these seem to be more frequent on Blue

08:30 Going to 1h30_45d field to run through focus for a couple of PA's. To ascertain that this defocus in X is not tilt related.

08:43 RB_rV superfoc at PA=0 and PA=180.

08:57 Running 1h30_45d field at SDTU and other filters (Blue: S, U-B and B-B and Red: z, Y, I-B) as a check on halo and focus issues - do these persist at this field which is not at low (>1.4) airmass?

09:00 Tried to analyze Blue focus sequence data. For PA=0, looks flat. For PA=180, I had some issues with obsutil.starfoc and it's "invalid flux profiles". Either this error, or a poor fit. I can repeat a PA=180 focus sequence after this.
  • PA=180 focus sequence.
  • dofpia
  • extra/intra focal images (RB_rVexin4.ob in /home/LBTO/Calib_OBs/FOCUS) 093532/093526 B/R
  • extra/intra at PA=180. (see more elongation in images).
  • Now run dofpia again... 09:42
  • RB_rV_superfoc_coarse ...
  • RB_rV_superfoc_coarse180...

09:56 dofpia... prior to observing 1h30_45d field at other filters whose focus offsets were tweaked: (Blue: r, g, U-B and Red: i-S, I-B and R-B)
  • see some elongation creeping in, so will obtain the coarse RB superfoc immediately afterwards.
  • dofpia after that. First Red correction is bigger than first blue correction (which was within conv limits already).

10:50 Loaded New collimation models. The ones I was using are renamed:

[telescope@obs2 PSF]$ cp SXPMLBCCollimation.dat SXPMLBCCollimation_inplace20110920.dat
[telescope@obs2 PSF]$ cp DXPMLBCCollimation.dat DXPMLBCCollimation_inplace20110920.dat
[telescope@obs2 PSF]$ cp DXPMLBCCollimation.20110918.dat DXPMLBCCollimation.dat
cp: overwrite `DXPMLBCCollimation.dat'? y
[telescope@obs2 PSF]$ cp SXPMLBCCollimation.20110918.dat SXPMLBCCollimation.dat
cp: overwrite `SXPMLBCCollimation.dat'? y

10:50 Steve clicked initialize & collimate. I will slew to zenith (85 deg). Since we're not at zenith, I run dofpia here, 85 deg.
  • At 85 deg elevation. dofpia failed on Blue because pupils near focus. Temperature correction is 0.683 for SX and 0.062 for DX. T=10C.
  • At 72 deg elevation, again the Blue are in focus. Should the temp corrections for SX and DX differ so much?
  • I don't immediately see anything crazy in the new collimation model temperature coeffs. But, for reference, the old collimation models show z temp corrections of 2.568 (SX) and 2.062 (DX).

11:01 Restoring old collimation models:
[telescope@obs2 PSF]$ cp SXPMLBCCollimation_inplace20110920.dat SXPMLBCCollimation.dat
cp: overwrite `SXPMLBCCollimation.dat'? y
[telescope@obs2 PSF]$ cp DXPMLBCCollimation_inplace20110920.dat DXPMLBCCollimation.dat
cp: overwrite `DXPMLBCCollimation.dat'? y

11:08 Slewing to N925 OSU monitor target.
  • at focus position. running dofpia.

11:26 Starting N925 OB which will observe at B, S, V on Blue and R on Red. On Red image 114458 (and probably the other ones too), we notice many dark blotches - size of stars, but some dust. Is this in all images, from last night too?

11:50 OB finished. Running dofpia (oops - forgot to offset to focus field and ran dofpia directly on galaxy field) and will run U-B observation on same object in case SDT_Uspec with halos is not useable.

11:58 Starting U-B OB on N925. Twilight has started but it looks dark outside. OB finished at 05:20 - still looked pretty dark but I could see dawn in E.

12:00 "allseeing" output looks good except for the images at the start of second half and a z-S image later. Going to take a 150s exposure at z on this field, to see if we get one good z-S image at least.

12:19 start z-S image on N925 (N925z). NOTE: FWHM on Red tech chip image is worse than on Blue tech chip image. This has been consistent tonight, through different filters. Saturated 150s image - too light. Take 10s image. 122829 for B. Focus on r-SLOAN B image does not look so good. This was one FFoffset changed the most...*will need to check r-SLOAN Blue focus offset*

12:32 Start flats
  • VR 4500/5400 in 1 sec. Take series of 2sec flats 10-20k
  • Bz 20k in 1 sec. Take series of 1sec flats. 29k (B filter hardware. 2011/09/20 12:41:32.580167 FILTERS WHEEL#2 timeout error on stop (1 ) [src/filters/filters.c:2245] ) Blue is stuck on 124120.
  • SY Was going to do SY but filter failure prevented this. Leave as is for further investigation...

12:45 Closing.

END OF NIGHT NOTES:
  • How can I run the tec chip FPIA offline, to see which stars it uses, rejects? Corrections were sent from guide/tech chips.
  • Note Red guide chip FWHM has been larger than Blue for almost all the night. I know Red guide chip is slightly out of focus, but usually FWHM is still smaller than Blue. Images do not seem to bear this out, but must review more carefully. (allseeing output attached).
  • Investigate Haloes on SDT_Uspec. Only on this filter, not on others.
  • Investigate Dark splotches on chip 1 Red (see 114458). Dust? Maybe these were there all along.
  • Chevron on LBC Red chip 2 still there.
  • Bias levels are approximately the following. These biases were NOT taken in controlled conditions - lights were on in dome and perhaps telescope was slewing during some. The few biases taken after 06UT had higher counts than those before and up to about 06UT.
chip Red Blue
1 420 (450) 500 (500)
2 450 (470) 493 (500)
3 385 (420) 490 (500)
4 510 (515) 495 (505)
  • There were some images, particularly at the start of the second half, which showed elongation at one edge and round images on the other. This seemed most prevalent for Blue. The guide star started to appear almost double. But we also saw this for Red. When the guide star was double, images on chip 1 appeared elongated up/down. See B/R 081113/081110 for a pair of B/R images which both show this non-uniform PSF pattern. (Note added later: This pair of images shows elongation consistent with rotator tracking problems on both Blue and Red, not the image quality problems discussed here) Object was setting - was collimation not keeping up? But focus corrections were being sent for the most part, although missed during U/S images. Maybe by the time V/B images were done, the focus had already deteriorated. At a higher field, this did not seem to be such a problem. *Data to assess image quality over the field were obtained. The images are listed in the file "iqdata" which is attached. These are:
    • series of superfoc sequences at PA=0 and PA=180.
    • extra/intra focal pair of images at PA=0 and PA=180.
    • series of coarse, delta=0.08mm, superfoc sequences at PA=0 and PA=180.
  • Check Blue r-SLOAN focus offset. This one was changed more than the others, from 0.07 to 0.04 mm. Only image taken at end of night and only 10-sec looks elongated.

-- OlgaKuhn - 20 Sep 2011

  • haloes around bright stars thru SDT_Uspec filter:
    SDTUhaloes.jpg

  • iqdata: Data to assess image quality over field

  • iqdata_app: More data to assess image quality, taken immediately after seeing elong images
I Attachment Action Size Date Who Comment
20110920.fwhmfwhm 20110920.fwhm manage 4 K 20 Sep 2011 - 13:52 OlgaKuhn Output of allseeing script
SDTUhaloes.jpgjpg SDTUhaloes.jpg manage 102 K 20 Sep 2011 - 09:10 OlgaKuhn haloes around bright stars thru SDT_Uspec filter
iqdataEXT iqdata manage 4 K 20 Sep 2011 - 14:06 OlgaKuhn Data to assess image quality over field
iqdata_appEXT iqdata_app manage 1 K 20 Sep 2011 - 14:45 OlgaKuhn More data to assess image quality, taken immediately after seeing elong images
Topic revision: r2 - 20 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback